4. Prove that the language L {07 1m0 lm, n 20 is not regular
Answer #1

Let p be the pumping length.

Let S = (0^p)(1^p)(0^p)

Breaking this into xyz.

As |xy|<=p and |y|>0 so y = 0^k for k>0

So x(y^2)z = (0^(p+k))(1^p)(0^p)